OpenStudy (angel_kitty12):

Can someone help me factor this polynomial? I'm very confused and have tried everything. (Algebra 2) 2x^3-98a^2x+28x^2+98x

OpenStudy (angel_kitty12):

\[2x^3-98a^2x+28x^2+98x\] Here is the equation

OpenStudy (loser66):

2 is in common , hence it is out x is in common,it is out also. You show me your work

OpenStudy (angel_kitty12):

okay i will try to do so. what i got was \[2x(x^2-49a^2+14x+49)\]

OpenStudy (angel_kitty12):

i got stuck here. I'm this can be simplified more and that this isn't the final answer... @Loser66

OpenStudy (angel_kitty12):

We just found the GCF but haven't factored it. Although, 14 is not factorable by 2

OpenStudy (angel_kitty12):

you cannot square 14

OpenStudy (loser66):

x^2+14x+49= (x+7)^2

OpenStudy (loser66):

See?

OpenStudy (angel_kitty12):

how did you get that?

OpenStudy (loser66):

\(x^2+2*7*x+7^2=(x+7)^2\)

OpenStudy (loser66):

Formula

OpenStudy (loser66):

If you don't like it, just go backward \((x+7)^2 =(x+7)*(x+7)=x^2+14x+49\)

OpenStudy (angel_kitty12):

I'm sorry if i am making this complicated but where did the' a' go? I'm usually very good at this but I just don't get this. I really appreciate you taking your time for me though.

OpenStudy (loser66):

Not yet. YOu must understand this step first

OpenStudy (angel_kitty12):

can you do it step by step for me?

OpenStudy (loser66):

Ok

OpenStudy (angel_kitty12):

I'm really trying to learn this for a test I have soon.

OpenStudy (loser66):

\[2x[(x+7)^2-49a^2]\] ok then?

OpenStudy (angel_kitty12):

okay that i get. you grouped the first three variable to get (x+7)^2

OpenStudy (loser66):

\(49a^2 = (7a)^2\) OK?

OpenStudy (angel_kitty12):

yes

OpenStudy (loser66):

so, \(2x[(x+7)^2-(7a)^2]\) ok?

OpenStudy (angel_kitty12):

yes

OpenStudy (loser66):

Do you know that \(A^2 - B^2 = (A+B)*(A-B)\)?

OpenStudy (loser66):